OpenStudy (anonymous):

the ratio of the corresponding edge lengths of two similar solids is 3:7 what is the ratio of there volumes?

OpenStudy (anonymous):

Typically the volume of an solid given an edge length \(x\) is going to be: \[ V = kx^3 \]I suppose it does depend a bit on the solid though.

OpenStudy (anonymous):

anyway, our new ratio is: \[ k(3)^3:k(7)^2=3^3:7^3 = 27:343 \]Hmm, but I need to think about this a bit more...